# An Analysis of Action-Value Temporal-Difference Methods That Learn State Values

### v1

@@ 4.2 Regularized Dueling Q-learning
We propose a new dueling algorithm which does not rely on subtracting the identifiability term in Eq. 8. We start with the most general decomposition of the action-value function,
@@
This decomposition is unidentifiable (Wang et al., 2016, Sec. 3); we can add a constant to $V(s)$ and subtract the same constant from $\operatorname{Adv}(s,a)$ without changing $Q(s,a)$. Consequently, there are infinitely many valid decompositions that satisfy Eq. 11 and no clear way for an algorithm to choose between them.
Dueling DQN seeks to accurately approximate both $v_{\pi}(s)$ and ${\operatorname{Adv}_{\pi}(s,a)\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}q_{\pi}(s,a)-v_{\pi}(s)}$ in order to estimate $q_{\pi}(s,a)$, which motivates the subtraction of the mean advantage to get Eq. 8 from Eq. 11. The mean advantage is meant to coarsely approximate the expected advantage under the target policy, but sacrifices the precise semantics of $v_{\pi}$ and $\operatorname{Adv}_{\pi}$. We propose an alternative approach in which we search for any two functions, $V(s)$ and $\operatorname{Adv}(s,a)$, that accurately reconstruct $q_{\pi}(s,a)$, further relaxing these semantics in order to expedite learning.
Let us consider the underdetermined system of equations in Eq. 11 to motivate our algorithm. For a particular state-action pair $(s,a)$, we can represent the specific solution found by Dueling Q-learning as an ordered pair: $(V(s),\operatorname{Adv}(s,a))$. This particular solution is just one of infinitely many ordered pairs that satisfy Eq. 11, which form a negatively sloped line in the $V$-$\operatorname{Adv}$ plane: $\operatorname{Adv}(s,a)=q_{\pi}(s,a)-V(s)$, depicted in Figure 4. When $Q(s,a)$ is initialized, the locus of valid initializations of $V(s)$ and $\operatorname{Adv}(s,a)$ is also a negatively sloped line: $\operatorname{Adv}(s,a)=Q(s,a)-V(s)$, again depicted in Figure 4. These two lines are parallel, and therefore the shortest directional path between them is the vector orthogonal to both of them—the minimum-norm projection. This fact remains true regardless of how $Q(s,a)$ is initialized, since the slope of both lines is always $-1$.
For this reason, we hypothesize that a good solution to Eq. 11 is the point on the solution line closest to the initialization of $V(s)$ and $\operatorname{Adv}(s,a)$, since it may require fewer updates to reach. In practice, because the state-value and advantage functions are typically initialized near zero for both tabular and deep RL, we can approximate this solution more simply as the closest point to the origin.
We therefore propose to apply $l_{2}$ regularization to Eq. 9 to simultaneously restore identifiability to Eq. 11 and encourage the algorithm to find this minimum-norm approximation. Specifically, we penalize the squared-error loss from Eq. 9 with a term of
@@
Since we now have $Q(s,a)=V(s)+\operatorname{Adv}(s,a)$ from Eq. 11 (no identifiability term), evaluating the chain rule in Eq. 9 gives us the following updates:
@@
where $\beta\in[0,1)$ is the regularization coefficient. We call this new algorithm Regularized Dueling Q-learning (RDQ). Specifically, we refer to this variant with the $l_{2}$ penalty as Soft RDQ. In practice, we found Soft RDQ to be noisy in the tabular setting, since the state and advantage values become leaky—they get pushed towards zero by the $1-\beta$ factor, and then have to compensate using the stochastic TD error, $\delta_{t}$. However, Soft RDQ is very useful in the case of function approximation, since the penalty in Eq. 12 is architecture-agnostic and can be easily combined with complex neural networks. We experiment with this in Section 4.3.
In the tabular setting, we do not need an $l_{2}$ penalty to obtain the desired minimum-norm solution. This is because differentiating the penalty in Eq. 12 with respect to $V(s)$ and then solving it for zero gives
@@
Furthermore, when we remove the $l_{2}$ penalty from Eqs. 13 and 14 by setting $\beta=0$, we see that both $V(s)$ and exactly one advantage value, $\operatorname{Adv}(s,a)$, are incremented by exactly the same amount at each time step—indicating that Eq. 15 is another invariant quantity. This implies that, regardless of how $V$ and $\operatorname{Adv}$ are initialized, Eq. 15 always remains true.444 That is, up to a fixed translation determined by the difference at initialization: $V_{0}(s)-\sum\limits_{a\in\mathcal{A}}\operatorname{Adv}_{0}(s,a)$. Therefore, rather than using an explicit $l_{2}$ penalty, we can directly apply the RDQ update without a penalty to achieve the minimum-norm solution in the tabular setting:
@@
We call this variant Hard RDQ, since it explicitly follows the minimum-norm path (on average) without the use of a soft penalty. Unfortunately, there does not appear to be a simple extension of this idea to the function approximation case. If we were to apply these update rules under function approximation, the different gradients calculated for $V(s)$ and $\operatorname{Adv}(s,a)$ would violate the invariant quantity in Eq. 15 and lose the minimum-norm convergence property.
To test this new algorithm, we repeat the off-policy control experiment from Section 3.2. We once again use Q-learning as a baseline, and then additionally compare Dueling Q-learning with Hard RDQ. The only minor modifications we make is that we now change the suboptimal reward from 00 to $-1$ (which does not change the optimal policy), set $\gamma=0.999$, and initialize $V(s)$ and $\operatorname{Adv}(s,a)$ using zero-mean Gaussian noise with a standard deviation of $2$. The rationale for the nonzero reward is so that $Q$ does not correctly estimate either of the actions’ true values initially (on average).
We plot the results in Figure 5. All experiment and plotting procedures remain the same as before. In contrast to the QVMAX algorithm tested earlier, both dueling methods dramatically outperform Q-learning, showing that the advantage decomposition is a highly effective strategy in general. In the largest instantiation of our MDP, where $\left\lvert\mathcal{A}\right\rvert=18$, Hard RDQ slightly outperforms Dueling Q-learning across the range of step sizes (see Figure 5; left, center). The spike at $\alpha=0.5$ is due to the fact that AV-learning methods sum the updates to $V$ and $\operatorname{Adv}$, making the total effective step size equal to $1$ and therefore unstable. Across the various MDP instances (see Figure 5; right), Hard RDQ performs slightly better than Dueling Q-learning in a small majority of cases, but both algorithms perform very well overall.
Figure 5: Off-policy control performance of Hard RDQ compared to Dueling Q-learning and Q-learning in the 4-state MDP. Experiment setup is the same as that of Figure 1.
@@ 4.3 Deep RL Experiments
Because Dueling Q-learning was originally proposed as a network architecture for Deep Q-Network (DQN; Mnih et al., 2015), we test the performance of RDQ in a deep RL setting, where a neural network is used to approximate the action-value function. Our benchmark is the MinAtar domain (Young & Tian, 2019), which includes five Atari-like games: Asterix, Breakout, Freeway, Seaquest, Space Invaders. The state representation for MinAtar is $10\times 10$ multi-channel binary images displaying various objects and their velocities. The reward function is the incremental game score.
@@
For RDQ and Dueling DQN, we create a dueling architecture following Wang et al.’s 2016 procedure. We duplicate the 128-unit hidden layer to branch the network in parallel streams, and then separately map these into linear outputs of size $\lvert\mathcal{A}\rvert$ and $1$ for estimating $\operatorname{Adv}(s,a;{\bm{\theta}}_{t})$ and $V(s,a;{\bm{\theta}}_{t})$, respectively. These are summed together (with broadcasting) to compute $Q(s,a;{\bm{\theta}}_{t})$. Dueling DQN additionally subtracts the identifiability term $\frac{1}{\lvert\mathcal{A}\rvert}\sum_{a^{\prime}\in\mathcal{A}}\operatorname{Adv}(s,a^{\prime};{\bm{\theta}}_{t})$ per Eq. 8.
@@
where ${\bm{\theta}}^{-}_{t}$ is the target network parameters copied from ${\bm{\theta}}_{t}$ every 1,000 time steps. The expectation in Eq. 16 is taken over the uniform distribution of samples $(S,A,R,S^{\prime})$ in $\mathcal{D}_{t}$; in practice, we approximate this with minibatches of size $32$. RDQ inherits the same loss, but adds the regularization term:
@@
Each agent was trained for a total of 10 million time steps. Every 1 million time steps, the agent was evaluated with an $\epsilon$-greedy policy for 1,000 episodes with $\epsilon=0.01$. In Figure 6, we plot the mean undiscounted return for the evaluation episodes as a function of training time; these results are averaged over 30 independent trials and the shading indicates 95% confidence intervals.
Figure 6 depicts our results. We see that between DQN and Dueling DQN, there does not appear to be a clearly superior algorithm, and one algorithm may outperform the other depending on the environment. However, we also see that RDQ significantly outperforms DQN and Dueling DQN in all five environments. Note that RDQ shares the same architecture as Dueling DQN. Although $\beta=10^{-3}$ works well in this setting, more experiments would be needed to determine its efficacy in other domains. Given that we did not tune $\beta$, it is also possible that performance could improve if tuned.
Figure 6: Deep RL results across five MinAtar games. Each algorithm is run for 30 seeds in each environment. The dark lines represent the mean across these 30 seeds, and the shaded region corresponds to a 95% confidence interval.
@@ 5 Conclusion
In this paper, we made several advances in the understanding of TD-learning algorithms in which state-value functions are learned in tandem with action-value functions. We showed that on-policy QV-learning algorithms have benefits for prediction, as in the original setting for which they were proposed, but tend to be much slower when the target policy is greedy. We also demonstrated that the prevailing off-policy control variant of QV-learning, QVMAX, is biased, and we introduced a new variant called BC-QVMAX which empirically restores convergence. Lastly, we introduced a novel Dueling Q-learning method called RDQ, which does not require the subtraction of an identifiability term, and has desirable properties in terms of the geometry of the learned value functions. In a deep RL setting based on the MinAtar games, we showed that RDQ greatly outperforms Dueling DQN, despite having identical network architectures and the same, well-tuned hyperparameters. Although there is still much to learn about these algorithms, our analysis helps to clarify their efficacy and distinguishing properties, and has already demonstrated potential for the development of new and effective RL algorithms.
@@ Appendix A Proofs
This section contains the proofs for all theoretical results in the paper. These results rely on the analysis of dynamic-programming operators for MDPs, which we introduce here rather than Section 2 for clarity of exposition.
@@
[…]
@@
We adopt the operator convention of Daley (2025), summarized in Table 1, which decomposes the standard Bellman operators into three fundamental operators: one for state transitions (${\bm{P}}$) and two for action selection (${\bm{E}}_{\pi}$ and $E$). The major benefit to this notation is that these fundamental operators are never overloaded; each applies to either ${\bm{v}}$ or ${\bm{q}}$ but not both. This allows us to expand the Bellman operators for ${\bm{v}}$ and ${\bm{q}}$ into unambiguous expressions below—especially useful for analyzing QV-learning methods.
@@
$T_{\pi}{\bm{q}}$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{r}}+\gamma{\bm{P}}{\bm{E}}_{\pi}{\bm{q}}\,,$
$T_{\pi}{\bm{v}}$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{E}}_{\pi}({\bm{r}}+\gamma{\bm{P}}{\bm{v}})\,.$
The Bellman expectation equations, Eqs. 1 and 4, can now be expressed succinctly as ${\bm{q}}_{\pi}=T_{\pi}{\bm{q}}_{\pi}$ and ${\bm{v}}_{\pi}=T_{\pi}{\bm{v}}_{\pi}$. In other words, the unique fixed point of $T_{\pi}$ is either ${\bm{q}}_{\pi}$ or ${\bm{v}}_{\pi}$ depending on whether it acts on action values or state values, respectively. Additionally, these fixed points are related by ${\bm{v}}_{\pi}={\bm{E}}_{\pi}{\bm{q}}_{\pi}$.
@@
$T{\bm{q}}$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{r}}+\gamma{\bm{P}}E{\bm{q}}\,,$
$T{\bm{v}}$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}E({\bm{r}}+\gamma{\bm{P}}{\bm{v}})\,.$
@@
The operator updates corresponding to the QV-learning, Eqs. 5 and 6, are
@@
The operator updates corresponding to Eqs. 5 and 7 are
@@
This mismatch stems from the fact that the updates to ${\bm{v}}$ are conditioned only on the states of the MDP and do not correct for the influence of taken actions. As such, the updates are subject to changes to the reward and next-state distributions induced by the particular behavior policy, $b$. This manifests as the discrepancy ${\bm{E}}_{b}\neq E$ in the above equations. These two operators only coincide in the serendipitous case that $b$ happens to be greedy with respect to the expected $1$-step returns, ${\bm{r}}+\gamma{\bm{P}}{\bm{v}}$, but not in general. Ultimately, the above analysis demonstrates that $\begin{bmatrix}{\bm{q}}_{*}\\ {\bm{v}}_{*}\end{bmatrix}$ cannot be the fixed point of QVMAX for an arbitrary behavior policy $b$ and MDP, which completes the proof. ∎
A.3 Proof of Proposition 3.3
@@
The operator updates corresponding to BC-QVMAX, Eqs. 5 and 10, are

### v2

@@ 4.2 Regularized Dueling Q-learning
We propose a new dueling algorithm which does not rely on subtracting the identifiability term in Eq.˜8. We start with the most general decomposition of the action-value function,
@@
This decomposition is unidentifiable (Wang et al., 2016, Sec. 3); we can add a constant to $V(s)$ and subtract the same constant from $\operatorname{Adv}(s,a)$ without changing $Q(s,a)$. Consequently, there are infinitely many valid decompositions that satisfy Eq.˜11 and no clear way for an algorithm to choose between them.
Dueling DQN seeks to accurately approximate both $v_{\pi}(s)$ and $\operatorname{Adv}_{\pi}(s,a)\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}q_{\pi}(s,a)-v_{\pi}(s)$} in order to estimate $q_{\pi}(s,a)$, which motivates the subtraction of the mean advantage to get Eq.˜8 from Eq.˜11. The mean advantage is meant to coarsely approximate the expected advantage under the target policy, but sacrifices the precise semantics of $v_{\pi}$ and $\operatorname{Adv}_{\pi}$. We propose an alternative approach in which we search for any two functions, $V(s)$ and $\operatorname{Adv}(s,a)$, that accurately reconstruct $q_{\pi}(s,a)$, further relaxing these semantics in order to expedite learning.
Let us consider the underdetermined system of equations in Eq.˜11 to motivate our algorithm. For a particular state-action pair $(s,a)$, we can represent the specific solution found by Dueling Q-learning as an ordered pair: $(V(s),\operatorname{Adv}(s,a))$. This particular solution is just one of infinitely many ordered pairs that satisfy Eq.˜11, which form a negatively sloped line in the $V$-$\operatorname{Adv}$ plane: $\operatorname{Adv}(s,a)=q_{\pi}(s,a)-V(s)$, depicted in Figure˜4. When $Q(s,a)$ is initialized, the locus of valid initializations of $V(s)$ and $\operatorname{Adv}(s,a)$ is also a negatively sloped line: $\operatorname{Adv}(s,a)=Q(s,a)-V(s)$, again depicted in Figure˜4. These two lines are parallel, and therefore the shortest directional path between them is the vector orthogonal to both of them—the minimum-norm projection. This fact remains true regardless of how $Q(s,a)$ is initialized, since the slope of both lines is always 1-1.
For this reason, we hypothesize that a good solution to Eq.˜11 is the point on the solution line closest to the initialization of $V(s)$ and $\operatorname{Adv}(s,a)$, since it may require fewer updates to reach. In practice, because the state-value and advantage functions are typically initialized near zero for both tabular and deep RL, we can approximate this solution more simply as the closest point to the origin.
We therefore propose to apply $l_{2}$ regularization to Eq.˜9 to simultaneously restore identifiability to Eq.˜11 and encourage the algorithm to find this minimum-norm approximation. Specifically, we penalize the squared-error loss from Eq.˜9 with a term of
@@
Since we now have $Q(s,a)=V(s)+\operatorname{Adv}(s,a)$ from Eq.˜11 (no identifiability term), evaluating the chain rule in Eq.˜9 gives us the following updates:
@@
where $\beta\in[0,1)$ is the regularization coefficient. We call this new algorithm Regularized Dueling Q-learning (RDQ). Specifically, we refer to this variant with the $l_{2}$ penalty as Soft RDQ. In practice, we found Soft RDQ to be noisy in the tabular setting, since the state and advantage values become leaky—they get pushed towards zero by the $1-\beta$ factor, and then have to compensate using the stochastic TD error, $\delta_{t}$. However, Soft RDQ is very useful in the case of function approximation, since the penalty in Eq.˜12 is architecture-agnostic and can be easily combined with complex neural networks. We experiment with this in Section˜4.3.
In the tabular setting, we do not need an $l_{2}$ penalty to obtain the desired minimum-norm solution. This is because substituting $\operatorname{Adv}(s,a)=Q(s,a)-V(s)$ into the penalty in Eq.˜12, differentiating with respect to $V(s)$, and setting the derivative to zero yields
@@
Furthermore, when we remove the $l_{2}$ penalty from Eqs.˜13 and 14 by setting $\beta=0$, we see that both $V(s)$ and exactly one advantage value, $\operatorname{Adv}(s,a)$, are incremented by exactly the same amount at each time step—indicating that Eq.˜15 is another invariant quantity. This implies that, regardless of how $V$ and $\operatorname{Adv}$ are initialized, Eq.˜15 always remains true.444 That is, up to a fixed translation determined by the difference at initialization: $V_{0}(s)-\!\!\sum\limits_{a\in\mathcal{A}}\operatorname{Adv}_{0}(s,a)$. Therefore, rather than using an explicit $l_{2}$ penalty, we can directly apply the RDQ update without a penalty to achieve the minimum-norm solution in the tabular setting:
@@
We call this variant Hard RDQ, since it explicitly follows the minimum-norm path (on average) without the use of a soft penalty. Unfortunately, there does not appear to be a simple extension of this idea to the function approximation case. If we were to apply these update rules under function approximation, the different gradients calculated for $V(s)$ and $\operatorname{Adv}(s,a)$ would violate the invariant quantity in Eq.˜15 and lose the minimum-norm convergence property.
To test this new algorithm, we repeat the off-policy control experiment from Section˜3.2. We once again use Q-learning as a baseline, and then additionally compare Dueling Q-learning with Hard RDQ. The only minor modifications we make is that we now change the suboptimal reward from 0 to 1-1 (which does not change the optimal policy), set $\gamma=0.999$, and initialize $V(s)$ and $\operatorname{Adv}(s,a)$ using zero-mean Gaussian noise with a standard deviation of $2$. The rationale for the nonzero reward is so that $Q$ does not correctly estimate either of the actions’ true values initially (on average).
We plot the results in Figure˜5. All experiment and plotting procedures remain the same as before. In contrast to the QVMAX algorithm tested earlier, both dueling methods dramatically outperform Q-learning, showing that the advantage decomposition is a highly effective strategy in general. In the largest instantiation of our MDP, where $\left\lvert\mathcal{A}\right\rvert=18$, Hard RDQ slightly outperforms Dueling Q-learning across the range of step sizes (see Figure˜5; left, center). The spike at $\alpha=0.5$ is due to the fact that AV-learning methods sum the updates to $V$ and $\operatorname{Adv}$, making the total effective step size equal to $1$ and therefore unstable. Across the various MDP instances (see Figure˜5; right), Hard RDQ performs slightly better than Dueling Q-learning in a small majority of cases, but both algorithms perform very well overall.
Figure 5: Off-policy control performance of Hard RDQ compared to Dueling Q-learning and Q-learning in the 4-state MDP. Experiment setup is the same as that of Figure˜1.
@@ 4.3 Deep RL Experiments
Because Dueling Q-learning was originally proposed as a network architecture for Deep Q-Network (DQN; Mnih et al., 2015), we test the performance of RDQ in a deep RL setting, where a neural network is used to approximate the action-value function. We implement our agents using the PFRL library (Fujita et al., 2021). Our benchmark is the MinAtar domain (Young & Tian, 2019), which includes five Atari-like games: Asterix, Breakout, Freeway, Seaquest, Space Invaders. The state representation for MinAtar is $10\times 10$ multi-channel binary images displaying various objects and their velocities. The reward function is the incremental game score.
@@
For RDQ and Dueling DQN, we create a dueling architecture following Wang et al.’s 2016 procedure. We duplicate the 128-unit hidden layer to branch the network in parallel streams, and then separately map these into linear outputs of size $\lvert\mathcal{A}\rvert$ and $1$ for estimating $\operatorname{Adv}(s,a;{\bm{\theta}}_{t})$ and $V(s;{\bm{\theta}}_{t})$, respectively. These are summed together (with broadcasting) to compute $Q(s,a;{\bm{\theta}}_{t})$. Dueling DQN additionally subtracts the identifiability term, $\frac{1}{\lvert\mathcal{A}\rvert}\sum_{a^{\prime}\in\mathcal{A}}\operatorname{Adv}(s,a^{\prime};{\bm{\theta}}_{t})$, per Eq.˜8.
@@
where ${\bm{\theta}}^{-}_{t}$ is the target network parameters copied from ${\bm{\theta}}_{t}$ every 1,000 time steps. The expectation in Eq.˜16 is taken over the uniform distribution of samples $(S,A,R,S^{\prime})$ in $\mathcal{D}_{t}$; in practice, we approximate this with minibatches of size $32$. RDQ inherits the same loss, but adds the regularization term:
@@
Each agent was trained for a total of 10 million time steps. Every 1 million time steps, the agent was evaluated with an $\epsilon$-greedy policy for 1,000 episodes with $\epsilon=0.01$. In Figure˜6, we plot the mean undiscounted return for the evaluation episodes as a function of training time; these results are averaged over 30 independent trials and the shading indicates 95% confidence intervals.
Figure˜6 depicts our results. We see that between DQN and Dueling DQN, there does not appear to be a clearly superior algorithm, and one algorithm may outperform the other depending on the environment. However, we also see that RDQ significantly outperforms DQN and Dueling DQN in all five environments. Note that RDQ shares the same architecture as Dueling DQN. Although $\beta=10^{-3}$ works well in this setting, more experiments would be needed to determine its efficacy in other domains. Given that we did not tune $\beta$, it is also possible that performance could improve if tuned.
Figure 6: Deep RL results for five MinAtar games. Averaged across 30 trials; shading represents 95% confidence intervals.
@@ 5 Conclusion
We made several advances in the understanding of TD algorithms in which state-value functions are learned in tandem with action-value functions. We showed that QV-learning algorithms have benefits for on-policy prediction—the original setting for which they were proposed—but tend to be much slower when the target policy is greedy. We also demonstrated that the prevailing off-policy control variant of QV-learning, QVMAX, is biased, and we introduced a new variant called BC-QVMAX which empirically restores convergence. Lastly, we introduced a novel Dueling Q-learning method called RDQ, which does not require the subtraction of an identifiability term, and has desirable properties in terms of the geometry of the learned value functions. In a deep RL setting based on the MinAtar games, we showed that RDQ greatly outperforms Dueling DQN, despite having identical network architectures and the same, well-tuned hyperparameters. Although there is still much to learn about these algorithms, our analysis helps to clarify their efficacy and distinguishing properties, and has already demonstrated potential for the development of new and effective RL algorithms.
@@ Appendix A Proofs
This section contains the proofs for all theoretical results in the paper. These results rely on the analysis of dynamic-programming operators for MDPs, which we introduce here rather than Section˜2 for clarity of exposition.
@@
$r(s,a)\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\sum_{r\in\mathcal{R}}\mathop{r}\sum_{s^{\prime}\in\mathcal{S}}\mathop{p(s^{\prime},r\mid s,a)}$
@@
We adopt the operator convention of Daley (2025), summarized in Table˜1, which decomposes the standard Bellman operators into three fundamental operators: one for state transitions ($\bm{P}$}) and two for action selection (${\bm{E}}_{\pi}$ and $E$). The major benefit to this notation is that these fundamental operators are never overloaded; each applies to either $\bm{v}$} or $\bm{q}$} but not both. This allows us to expand the Bellman operators for $\bm{v}$} and $\bm{q}$} into unambiguous expressions below—especially useful for analyzing QV-learning methods.
@@
$T_{\pi}{\bm{q}}$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{r}}+\gamma{\bm{P}}{\bm{E}}_{\pi}{\bm{q}}\,$,
$T_{\pi}{\bm{v}}$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{E}}_{\pi}({\bm{r}}+\gamma{\bm{P}}{\bm{v}})\,.$
The Bellman expectation equations, Eqs.˜1 and 4, can now be expressed succinctly as ${\bm{q}}_{\pi}=T_{\pi}{\bm{q}}_{\pi}$ and ${\bm{v}}_{\pi}=T_{\pi}{\bm{v}}_{\pi}$. In other words, the unique fixed point of $T_{\pi}$ is either ${\bm{q}}_{\pi}$ or ${\bm{v}}_{\pi}$ depending on whether it acts on action values or state values, respectively. Additionally, these fixed points are related by ${\bm{v}}_{\pi}={\bm{E}}_{\pi}{\bm{q}}_{\pi}$.
@@
$T{\bm{q}}$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\bm{r}}+\gamma{\bm{P}}E{\bm{q}}\,$,
$T{\bm{v}}$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}E({\bm{r}}+\gamma{\bm{P}}{\bm{v}})\,.$
@@
The operator updates corresponding to the QV-learning, Eqs.˜5 and 6, are
@@
The operator updates corresponding to Eqs.˜5 and 7 are
@@
This mismatch stems from the fact that the updates to $\bm{v}$} are conditioned only on the states of the MDP and do not correct for the influence of taken actions. As such, the updates are subject to changes to the reward and next-state distributions induced by the particular behavior policy, $b$. This manifests as the discrepancy ${\bm{E}}_{b}\neq E$ in the above equations. These two operators only coincide in the serendipitous case that $b$ happens to simultaneously maximize the expected $1$-step returns, ${\bm{r}}+\gamma{\bm{P}}{\bm{v}}$, but not in general. Ultimately, the above analysis demonstrates that $\begin{bmatrix}{\bm{q}}_{*}\\ {\bm{v}}_{*}\end{bmatrix}$ cannot be the fixed point of QVMAX for an arbitrary behavior policy and MDP, which completes the proof. ∎
A.3 Proof of Proposition˜3.3
@@
The operator updates corresponding to BC-QVMAX, Eqs.˜5 and 10, are